Multiple Choice Questions

  1. A company makes and sells product A and B. Twice as many units of product B are made and sold as that of product A. Each unit of product A makes a contribution of $10 and each unit of product B makes a contribution of $4. Fixed costs are $90,000.

    What is the total number of units which must be made and sold to make a profit of $45,000?

    A 7500 B 22,500 C 15,000 D 16875

  2. B Ltd is considering changing the way it is structured by asking its employed staff to become freelance. Employees are currently paid a fixed salary of $240,000 per annum, but would instead be paid $200 per working day. On a typical working day, staff can produce 40 units. Other fixed costs are $400,000 pa.

    The selling price of a unit is $60 and material costs are $20 per unit.

    What will be the effect of the change on the breakeven point of the business and the level of operating risk?

    1. A The breakeven point reduces by 6,000 units and the operating risk goes down.

    2. B The breakeven point reduces by 4,571 units and the operating risk goes down

    3. C The breakeven point reduces by 4,571 units and the operating risk goes up.

    4. D The breakeven point reduces b 6,000 units and the operating risk goes up.

  3. The CS ratio for a business is 0.4 and its fixed costs are $1,600,000. Budget revenue has been set at 6 times the amount of the fixed costs.

    What is the margin of safety in % measured in terms of revenue?

  4. A company manufactures and sells a single product with a variable cost per unit of $36. It has a contribution ratio of 25%. The company has weekly fixed costs of $18,000.

    What is the weekly breakeven point, in units? A 1,500
    B 1,600
    C 1,800

    D 2,000

1

2

5. The management accountant of C plc has calculated the firm’s breakeven point from the following data:

Selling price per unit $20 Variable costs per unit $8 Fixed overheads for next year $79,104

It is now expected that the product’s selling price and variable cost will increase by 8% and 5.2% respectively.

These changes will cause C plc’s breakeven point for next year to:

  1. A Rise by 9.0%

  2. B Rise by 2.8%

  3. C Fall by 2.8%

  4. D Fall by 9%
Answer #1

1. Assume the number of units produced by product A is "X"

So the number units produced by product B is "2X"

Profit = Contribution by each product * number of units of production of Each Product - Fixed Costs

In the given case :

Fixed cost =90,000

Profit =45000

Contribution for Product A =10

Contribution for Product B =4

So as per equatipon

45000=X*10+2X*4-90000

45000=18X-90000

18X=45000+90000=135000

X=135000/18=7500

So Product A units Is 7500 units

Product b units is 7500*2 =15000units

SO total units =7500+15000=22500 units

The total number of units which must be made and sold to make a profit of $45,000 is 22500 units

So answer is B 22500

2.1st we have to find of break even point before change

BEP Units = Fixed Cost/(selling price per unit- variable cost)

Fixed cost =240000+400000=640000

Selling price per unit =60

Variable cost = 20

BEP units = 640000/(60-20)=16000 units

SO existing BEP =16000 units

Introduction of new policy by paying 200 per working for 40Units production per day :

so per unit staff variable cost =200/40=5

Now fixed cost is 400000 only

Now Variable cost is 25

After change BEP =400000/(60-25)=11429 units

Break even Points reduces =16000-11429=4571 units

So Answer is  B The breakeven point reduces by 4,571 units and the operating risk goes down

3.Here is the Information

CS(Contribution to Sales) Ratio =0.4 =Contribution/Sales

Fixed Cost=1600000

Budgeted Revenue = 6 times of Fixed cost=6*1600000=9600000

Margin of Safety in Revenue terms =Sales-Break Even Sales

Margin of Safety Ratio =Margin of Safety/Sales*100

From the above

Contribution to Sales Ration =0.4=Contribution/9600000

Contribution = 9600000*.4=3840000

Variable Cost = 9600000-3840000=5760000

Variable cost percentage on Sales is 60%

Breaken Point =Fixed Cost/Contribution%

=1600000/.4=4000000

Break Even Sales(Revenue)=4000000

Actual Sales =9600000

Margin of Safety Sales = 9600000-4000000=5600000

Margin of Safety Ratio = 5600000/9600000*100=58.33%

Margin of safety is 58.33% measured in terms of revenue

4. BEP Units = Fixed Cost/(selling price per unit- variable cost)

Fixed cost =18000

Variable cost per unit = 36

Contribution Ration =25% =Contribution Per Unit/Selling Price per Unit

Selling price per unit = Varible Cost/(100%-25%)=36/75%=48

BEP Units = 18000/(48-36)=1500 units

Answer : The weekly breakeven point, in units is  A 1,500

5.

1st we have to find of break even point before change

BEP Units = Fixed Cost/(selling price per unit- variable cost)

Fixed cost =79104

Selling price per unit =20

Variable cost = 8

BEP units = 79104/(20-8)=6592 units

SO existing BEP =6592 units

the product’s selling price and variable cost will increase by 8% and 5.2% respectively.

Now Selling price per unit =20*1.08=21.60

NOw Variable cost per unit =8*1.052=8.416

After change BEP =79104/(21.60-8.416)=6000 units

Fall in Break even for next year from 6592 to 6000 units

So % of Fall in Break even Poins =(6592-6000)/6592=592/6592=8.98% =9%

Answer : D Fall by 9%